solve using substitution method
x - 5y = 17
-5x - 6y = 8

Answer by mananth(16946)

x-5y=17 .............1
6y=-8-5x ---------2
/6=
y=-1.33 -0.83x
Plug the value of y in (1)
x-5(-1.33-0.83x) =17
x+6.67+ 4.17x=17
x+4.17x= 17-6.67
5.17x=10.33
/5.17
x=2
Plug the value of x in (1)
x-5y=17
1*2-5y=17
2-5y=17
-5y=15
/-5
y=-3